The measure of one acute angle of a right triangle is 12 degrees more than the other acute angle. What is the measure of the smallest acute angle?
Answer by Theo(13342)

x = smaller angle
x + 12 = larger angle.
sume of the angles is equal to 90 degrees since this is a right triangle.
x + x + 12 = 90
combine like terms to get:
2x + 12 = 90
subtract 12 from both sides to get:
2x = 78
divide both sides by 2 to get:
x = 39
that's the smaller angle.
x+12 = 51
that's the bigger angle.
the sum of the angles is 39 + 51 = 90 which is good.